OpenStudy (anonymous):

Solve the equation sin(x) = 1/2 on the interval x E [0, pi]. (Angles are measured in radians.)

OpenStudy (anonymous):

\[\frac{\pi}{6}\] and \[\frac{5\pi}{6}\]

OpenStudy (anonymous):

tyler can you lay out the steps pls

OpenStudy (anonymous):

All i did was look on the unit circle where sin was positive 1/2.. Sine is positive in the first 2 quadrants (which is what your interval is)